Hodgkiss Corporation is evaluating an extra dividend versus a share repurchase. In either case, $21,060 would be spent. Current earnings are $3.60 per share, and the stock currently sells for $90 per share. There are 3,900 shares outstanding. Ignore taxes and other imperfections.

What will the company’s EPS and PE ratio be under the two different scenarios? (Do not round intermediate calculations and round your answers to 2 decimal places, e.g., 32.16.)

Extra Dividend Share Repurchase
EPS $ $
PE Ratio
Answer #1

If 21,060 would be spent each shareholder will receive 21,060/3900 = 5.4$ per share as dividend and price will reduce by 5.4

EPS will be same 3.6$ per share

PE ratio will be - price/EPS = (90 - 5.4)/3.6
= 23.5

Or

And 21,060/90 = 234 shares will be purchased.

Total earnings = 3.6*3900
= 14,040$

New EPS = 14,040/(3900-234)
= 3.83$

New PE ratio = 90/3.83
= 23.5